OpenStudy (anonymous):

how do i find the derivative of f(x) = log base 10 (2x^3 - 4x) + ln(sin^2x)?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

\[\frac{d}{dx}\log_b(f(x))=\frac{f'(x)}{\ln(b) f(x)}\]

OpenStudy (anonymous):

and \[\frac{d}{dx}\ln(f(x))=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

what will the answer be?...

OpenStudy (lalaly):

Try to find it ..

OpenStudy (anonymous):

To start, split it into the two separate parts: \( \log_{10}(2x^3-4x) \) and \(\ln(\sin^2x)\), then derive those parts each using the two formulas that satellite gave you.

OpenStudy (anonymous):

i already did the problem am checkin to see if i got the right answer.... i got 2 cot x + (4sec^2/tan x) - 4x/(x^2 + 1)

OpenStudy (anonymous):

That's not at all what I got. Can you show your work?

OpenStudy (anonymous):

The \(2\cot x\) part is right, but I don't know where you got the rest from.

OpenStudy (anonymous):

my teacher marked it correct..

OpenStudy (anonymous):

can you show your workout? thank

OpenStudy (lalaly):

\[\frac{d}{dx}(\log_{10}(2x^3-4x)+\ln(\sin^2x))\]\[=\frac{6x^2-4}{\ln(10)(2x^3-4x)}+\frac{2cosxsinx}{\sin^2x}\]\[=\frac{2(3x^2-2)}{2\ln(10)(x^3-2)}+2cotx\]

OpenStudy (anonymous):

thanks this looks close to what i got..